One-Year Trailing Returns
Miranda Fund S&P 500
Return 10.5 % −20.8 %
Standard deviation 36.0 % 41 %
Beta 1.30 1.00

For the entire year her asset class exposures averaged 50% in stocks and 50% in cash. The S&P’s allocation between stocks and cash during the period was a constant 94% and 6%, respectively. The risk-free rate of return was 3%.

What is the M2 measure for Miranda? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Return - 10.5%

Standard Deviation - 36%

Beta -1.3

Sharpe Ratio - .2083
Answer #1

M2 = Modigliani measure of risk adjusted performance

M2 = Sharpe ratio*Std Deviation of Market + Risk Free Return

= 0.2083*41+3.5 = 8.5403+3.5 = 12.0403